A sample of xenon gas occupies a volume of 7.06 L at 445 K. If the pressure remains constant, at what temperature will this same xenon gas sample have a volume of 3.88 L?
Answer #1

V1/T1 = v2/T2

V1 = 7.06 lt

T1 = 445 k

V2 = 3.88lt

T2 = ?

7.06/445 = 3.88/T2

T2 = 244.56 K